Circular No. 8542

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ck Independent discoveries of a supernova in the Abell galaxy cluster

1656 have been reported on unfiltered CCD images by H. Pugh and W. Li (LOSS/KAIT; cf. IAUC 8541) and by R. Quimby, F. Castro, P. Hoeflich, J. C. Wheeler (all at the University of Texas), and C. Gerardy (of Imperial College); Quimby’s group used the ROTSE-IIIb telescope (cf. IAUC 8508). Pugh and Li provide the following precise position for SN 2005ck: α = 13^{h}02^{m}18^{s}.72, δ = +28^{o}20′45′′.5 (equinox 2000.0), which is 58′′.3 east and 24′′.3 south of the center of an apparent host galaxy. Quimby et al. report position end figures 18^{s}.77, 43′′.8 for the new object. Approximate magnitudes for SN 2005ck: 2004 Dec. 15, [18.8 (ROTSE-IIIb); 2005 Jan. 14, [18.8 (ROTSE-IIIb); Apr. 17.26 UT, [19.5 (KAIT); May 23.25, [18.5 (KAIT); June 1.26, 19.0: (KAIT; hint of object near limit of image); 5.27, 18.7 (ROTSE-IIIb); 8.25, 18.6 (ROTSE-IIIb); 12.24, 18.6 (KAIT); 13.24, 18.5 (KAIT). Quimby adds that a spectrum (range 420–890 nm) of SN 2005ck, obtained on June 13.22 with the 9.2-m Hobby-Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by S. C. Odewhan and E. Terrazas, shows it to be a type-Ia supernova; the spectrum is very similar to that of SN 1994D near maximum light (Patat et al. 1996, MNRAS 278, 111). Using 1994D as a template, they find an approximate redshift of z = 0.08, ruling out any association to the neighboring Coma-cluster galaxies, leaving the host as yet unidentified.

SUPERNOVA 2005ci IN NGC 5682 M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrogram (range 340–740 nm) of SN 2005ci (cf. IAUC 8541), obtained by M. Calkins on June 12.34 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-II supernova. The spectrum consists of a flat continuum and P-Cyg lines of Hα and Hβ. Adopting the NED recession velocity of 2291 km/s for the host galaxy (from Falco et al. 1999, PASP 111, 438), the expansion velocity derived from the minimum of the Hβ line is ∼ 13000 km/s.

VARIABLE STAR IN NORMA Further to IAUC 8540, S. Laloe (Centre de Donnees de Strasbourg)

writes that the so-called USNO-A1.0 star of red mag 18.4 is actually a USNO-B1.0 star.

2005 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green